OpenStudy (anonymous):

Alrite here's another. The length of a rectangle is twice its width. If the area of the rectangle is 50ft , find its perimeter. r

OpenStudy (anonymous):

ok so this is a system of equations: lets say: x=width y=length so the wuestion tells us that: y=2x and it says that: 50=(y)(x) if we substitute y into x: 50=(2x)(x) 25=x^2 x=5 using our equations: y=2(5)=10 so the perimeter of the rectangle is just: P=2x+2y P=2(5)+2(10)=30